# Canonical scattering problem in topological metamaterials: Valley-Hall modes through a bend

### v1

@@ IV.3 Influence of the ribbon width
Figure 15: Influence of the ribbon width $N$ on the reflection coefficient for fixed onsite potential $u=0.3$.
@@ IV.4 Influence of the onsite potential
Figure 16: Influence of the onsite potential $u$ on the reflection coefficient. The width of the ribbon is fixed to $N=12$ for the bridge interface and $N=13$ for the zigzag interface.
V Extension to other interface types
In the previous sections, we have constructed and used the transfer matrix formalism to quantitatively analyse the scattering of valley Hall modes through sharp bends in graphene ribbons with zigzag edges. The construction of the propagating and evanescent modes via the transfer matrix presented in Sec. III.1 was specific to this particular ribbon configuration with zigzag edges. We show here that it is possible to generalise the construction such that it can be applied to arbitrary ribbons. As an illustration, we then apply it to study the scattering of interface modes in graphene ribbons with armchair edges.
V.1 Transfer matrix for arbitrary ribbons
As for the bend cell, we write the eigenvalue equation as
$\bm{M}_{1}\Psi_{n+1}=\bm{M}_{2}\Psi_{n}.$ (82)
$\Psi_{n+1}=\bm{M}\Psi_{n},$ (83)
where
V.1.1 Connection with graphene ribbons with zigzag edges
$\text{det}\left[E^{2}-(\bm{1}+\bm{J}^{T})(\bm{1}+\bm{J})\right]\neq 0.$ (87)
Therefore, the spectrum of $\bm{H}_{0}$ is given by the condition
$\bm{\mathcal{G}}_{aa}$ $=$ $E\left[E^{2}-\bm{Q}\bm{Q}^{T}\right]^{-1},$ (90)
$\bm{\mathcal{G}}_{bb}$ $=$ $E\left[E^{2}-\bm{Q}\bm{Q}^{T}\right]^{-1},$ (92)
V.2 Application to ribbons with armchair edges
V.2.1 Graphene ribbons
$p_{n}=\frac{n}{N+1}\pi,\quad\text{with}\quad n=1,2,...,N.$ (94)
The energy bands are then given by
V.2.2 Ribbons with an interface
VI Conclusion
@@
Our observations reinforce the idea that the high transmission of valley Hall modes across bends is better explained in terms of a good mode matching on each side of the bend rather than in terms of the topological protection of the valley index since only the $32\pi/$ 3 bends allows the valley conservation.
Finally, while the main part of the study was devoted to the propagation of valley Hall modes, we have shown in Sec. V that the tools and methods applied to ribbons with zigzag edges can be generalized to characterise the propagation in arbitrary ribbons. As an illustration, we have described and computed the reflection and transmission coefficients of interface modes in ribbons with armchair edges through a bend at a $\pi/3$ angle. We expect that our method can be used to further characterise mode propagation in a wide range of discrete lattices. The investigation of scattering scenarios in tight-bindings models with different geometries through the efficient and accurate transfer matrix formalism may then guide the design of future continuous phononics and photonics crystals and other metamaterials.
@@ Acknowledgement
This work received support from the French government under the France 2030 investment plan, as part of the Initiative d’Excellence d’Aix-Marseille Université - AMIDEX - AMX-19-IET- 010.

### v2

@@ IV.3 Influence of the ribbon width
Figure 15: Influence of the ribbon width $N$ on the reflection coefficient for fixed onsite potential $u=0.3$. For this value of the onsite potential, we have $N_{c}\approx 3.4$.
@@ IV.4 Influence of the onsite potential
Figure 16: Influence of the onsite potential $u$ on the reflection coefficient. The width of the ribbon is fixed to $N=12$ for the bridge interface and $N=13$ for the zigzag interface. For these ribbon widths, we have $u_{c}\approx 0.08$.
Overall, we note that the dependence of the reflection coefficient on the external parameters, $u$ and $N$, does not display a monotonous nor simple behavior. In particular, even though the modes are localised near the interface, the localisation length depends on those external parameters. To estimate this, a site localisation number $N_{\rm loc}=-\ln(\lambda_{-})^{-1}$ can be defined, where $\lambda_{-}$ is the eigenvalue given in Eq. (98). When the width of the ribbon is comparable to $N_{\rm loc}$, finite size effects are expected to play a significant role. Using the results detailed in Appendix B, we have that $N_{\rm loc}\geq N_{c}=\ln(u-\sqrt{1+u^{2}})^{-1}$, which implies that $N_{\rm loc}$ goes to infinity when $u$ goes to 0. Equivalently, we can define a critical onsite potential $u_{c}$ as a function of $N$ by inverting the previous relation. Since the mode profiles are intrinsically different whether $N$ is larger or smaller than $N_{c}$ (equivalently with $u$ and $u_{c}$), it is natural to expect different behaviours in the two regimes. For example, when $u=0.3$ as in Fig. 15, we have $N_{c}\approx 3.4$.
V Conclusion
@@
Our observations reinforce the idea that valley conservation is not correlated with a high transmission of interface modes across bends. In fact, in some configuration, high transmission occurs in spite of a change of the valley index. This is true for the case of a sharp turn at an angle $\pi/3$ as investigated in this study or in the case of a straight ribbon where the onsite potential is suddenly flipped, as shown in the supplementary material Sec. C. Indeed, we have observed high transmission of interface modes across $\pi/3$ bends, a configuration which connects different valleys on each side of the defect. To further support this idea, we have compared in the supplemental material Sec. D, the scattering coefficients of valley Hall modes to those of a bulk mode of a graphene ribbon, and it was observed that both the interface and the bulk modes have similar reflection and transmission coefficients in that specific geometrical configuration.
Finally, while the main part of the study was devoted to the propagation of valley Hall modes, we show in the supplementary material Sec. E that the tools and methods applied to ribbons with zigzag edges can be generalized to characterise the propagation in arbitrary ribbons. As an illustration, we have described and computed the reflection and transmission coefficients of interface modes in ribbons with armchair edges through a bend at a $\pi/3$ angle. We expect that our method can be used to further characterise mode propagation in a wide range of discrete lattices. The investigation of scattering scenarios in tight-bindings models with different geometries through the efficient and accurate transfer matrix formalism may then guide the design of future continuous phononics and photonics crystals and other metamaterials.
A particularly interesting future application of our methods would be the quantitative study of backscattering in analogue QSH systems. As previously mentioned, analogue QSH devices strongly rely on spatial symmetries to allow for the existence of topological protected modes. These spatial symmetries are inevitably broken when considering finite size systems and defects. For this reason, and because analogue QSH systems are reciprocal, we do not expect their edge modes to be perfectly immune to backscattering. Applying our methods to QSH systems would allow for a better characterisation of the scattering properties of edge modes in such devices and in classical topological insulators more generally.
@@ Acknowledgement
This work received support from the French government under the France 2030 investment plan, as part of the Initiative d’Excellence d’Aix-Marseille Université - A*MIDEX - AMX-19-IET- 010.
@@ B.2 Interface states of the periodic chain
Appendix C An example where valley change is unavoidable
We consider here a simple configuration where the valley index necessarily changes during the scattering process. The configuration consists in a straight ribbon with an interface (shown in red line in Fig. 20) where the onsite potential $u>0$ is suddenly switched to $-u$. This abrupt change of the onsite potential happens at the blue interface in Fig. 20.
Figure 20: Scattering in a straight ribbon with flip of the onsite potential. Panel a) shows a particular scattering configuration for a ribbon with $N=12$ and $u=0.4$ at energy $E=0$, together with the dispersion relation on each side of the defect and the mode profiles at $E=0$. The arrows represent the direction of propagation of the modes and their color indicates their valley index. Panel b) shows the defect cell, the red and blue line corresponds to the interfaces where the onsite potential changes sign. Panel c) shows the reflection and transmission coefficients spectrum for the entire energy range where the valley Hall modes are present.
In Fig. 20-a), we show the dispersion relation on both side of the defect, and we can directly see that the spectrum of the two ribbons are opposite. This implies in particular that if the right-moving valley Hall mode in the ribbon with $u>0$ has momentum $k$, then it will have momentum $-k$ in the ribbon with $u<0$, i.e. it will be located in the opposite valley. In addition, right-moving and left-moving valley Hall modes are located, by construction in opposite valleys. Hence, any scattering in such configuration will consist in an incoming mode in a specific valley, being either transmitted or reflected to modes with an opposite valley index. Therefore, a change in the valley index is unavoidable and there is a priori no way of distinguishing the preferred scattering channel between the transmitted and the reflected mode.
Further insights can nonetheless be gained by inspecting the modes on each side of the defect. As shown in the supplemental material Sec. B, interface modes will have symmetric or anti-symmetric profiles, depending on the type of interface and the sign of the onsite potential. This is illustrated in Fig. 20-a) where the right-moving modes are shown on either side of the ribbon. We can see in particular that to the left of the defect, the valley Hall modes are symmetric across the red interface while they are anti-symmetric to the right of the defect. Therefore, scattering from one side to the defect to the other will require not only to change the valley index but also to change the symmetry of the modes. Since the defect is itself mirror symmetric across the horizontal interface, see Fig. 20-b), it cannot mix the symmetric and anti-symmetric components of the modes. Therefore, we can anticipate, from symmetry consideration, that a incoming mode will be fully reflected at the switching interface. This is indeed what is observed in Fig. 20-a) and c).
Appendix D Trivial vs topological modes
One of the main results of this study is the observation of a high transmission of interface modes across defects independently of the conservation of the valley index. This observation suggests that the topological nature of the modes involved does not determine the amount of backscattering.
Figure 21: Comparison between interface and bulk modes scattering in a sharp bend at angle $32\pi/$ 3 in a ribbon with $N=5$ and $u=0.4$. Panel a) shows the dispersion relation of the ribbon, the green region shows part of the energy range where interface modes are present and the orange region shows the region where a single pair of bulk modes are propagating. Panel b) shows the scattering coefficients for both interface and bulk modes, we can see that the amount of backscattering is similar for the two types of modes.
Here, we compare the scattering coefficients of a interface valley Hall mode with these of a trivial bulk mode of the ribbon, in the case of a sharp turn at angle $32\pi/$ 3 for a ribbon with $N=5$ and $u=0.4$. The dispersion relation for the ribbon and the scattering coefficients are shown in Fig. 21. Here we have chosen a relatively narrow ribbon to limit the number of propagating modes. This allows us to define an energy range where a single pair of propagating bulk modes is present, shown in orange. This energy range is comparable to the energy range shown in green where a single pair of propagating interface modes is present. We can see in Fig. 21-b) that the scattering coefficients are similar for both interface and bulk modes, reinforcing our observation that the amount of backscattering is not characterised by topological properties.
Appendix E Extension to other interface types
In the main part of the text, we have constructed and used the transfer matrix formalism to quantitatively analyse the scattering of valley Hall modes through sharp bends in graphene ribbons with zigzag edges. The construction of the propagating and evanescent modes via the transfer matrix presented in Sec. III.1 was specific to this particular ribbon configuration with zigzag edges. We show here that it is possible to generalise the construction such that it can be applied to arbitrary ribbons. As an illustration, we then apply it to study the scattering of interface modes in graphene ribbons with armchair edges.
E.1 Transfer matrix for arbitrary ribbons
As for the bend cell, we write the eigenvalue equation as
$\bm{M}_{1}\Psi_{n+1}=\bm{M}_{2}\Psi_{n}.$ (106)
$\Psi_{n+1}=\bm{M}\Psi_{n},$ (107)
where
E.1.1 Connection with graphene ribbons with zigzag edges
$\text{det}\left[E^{2}-(\bm{1}+\bm{J}^{T})(\bm{1}+\bm{J})\right]\neq 0.$ (111)
Therefore, the spectrum of $\bm{H}_{0}$ is given by the condition
$\bm{\mathcal{G}}_{aa}$ $=$ $E\left[E^{2}-\bm{Q}\bm{Q}^{T}\right]^{-1},$ (114)
$\bm{\mathcal{G}}_{bb}$ $=$ $E\left[E^{2}-\bm{Q}\bm{Q}^{T}\right]^{-1},$ (116)
E.2 Application to ribbons with armchair edges
E.2.1 Graphene ribbons
$p_{n}=\frac{n}{N+1}\pi,\quad\text{with}\quad n=1,2,...,N.$ (118)
The energy bands are then given by
E.2.2 Ribbons with an interface